NITROGEN-BROADENED ^{13}CH_{4} AT 80 TO 296 K

\underline{M. A. H. SMITH}, Science Directorate, NASA Langley Research Center, Hampton, VA 23681; K. SUNG, L. R. BROWN, T. J. CRAWFORD, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr.,Pasadena, CA 91109; A. W. MANTZ, Dept. of Physics, Astronomy and Geophysics, Connecticut College, New London, CT 06320; V. MALATHY DEVI and D. CHRIS BENNER, The College of William and Mary, Williamsburg, VA 23187.

High-resolution spectra of the ν_{4} fundamental band of ^{13}CH_{4} broadened by N_{2} at temperatures relevant to the atmosphere of Titan were recorded using temperature-controlled absorption cells^{a} installed in the sample compartment of a Bruker IFS-125HR Fourier Transform spectrometer (FTS) at the Jet Propulsion Laboratory (JPL). Analysis of these spectra using multispectrum fitting^{b} has determined half widths, pressure-induced shifts, line mixing parameters and their temperature dependences for transitions belonging to a number of P- and R-branch J-manifolds. The analysis examined in detail the temperature-dependence of N_{2}-broadened half width and pressure-induced shift coefficients over the entire temperature range from 80 K to 296 K. The results are compared with other published measurements of N_{2}- and air-broadened methane parameters.^{c}

aK. Sung, A. W. Mantz, M. A. H. Smith, L. R. Brown, T. J. Crawford, V. Malathy Devi and D. C. Benner, JMS \underline{262} (2010)

122-134.

bD. C. Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. A. Atkins, JQSRT \underline{53} (1995) 705-721.

cResearch described in this paper was performed at Connecticut College, the College of William and Mary, NASA Langley Research Center and the Jet Propulsion Laboratory, California Institute of Technology, under contracts and cooperative agreements with the National Aeronautics and Space Administration.